Bailand Company purchased a building for $366,000 that had an estimated residual value of $16,000 and an estimated service liASSETS REVENUE 111 Cash 411 Sales Revenue 121 Accounts Receivable 141 Inventory EXPENSES 500 Cost of Goods Sold 181 BuildingGeneral Journal Bailand estimates that the asset has 8 years life remaining (for a total of 12 years). Prepare the journal ePAGE 16 GENERAL JOURNAL DATE ACCOUNT TITLE POST. REF. DEBIT CREDIT Prepare the journal entries on December 31 to record the pPAGE 16 GENERAL JOURNAL DATE ACCOUNT TITLE POST. REF. DEBIT CREDIT
Answer #1

Building Cost                $366,000

Residual Value              $16,000

Estimated Service Life 10 years

Purchased            4 years ago

1. Balland estimates that the asset has 8 years’ life remaining (for a total of 12 years).

Straight Line Method

Year

Book Value of Asset at Beg. Of Year

Depreciation Exp

Accumulated Depreciation at Year End

Book Value at Year End(cost-Accudepryear end)

1-4

366,000

35,000

140,000

226,000

5

226,000

26,250

166,250

199,750

Straight-Line Depreciation Expense = Cost – Estimated Residual

Estimated Service Life

SL Dep Expe = $366,000 - $16,000/10 years = $35,000/year

Accumulated depreciation at the end of 4 years = $35,000 x 4 = $140,000

Book Value at the begging of year 5 = 366,000-140,000 = 226,000

Depreciation of the rest of the years

Depreciation = Remaining Book Value – Residual Value                226,000 – 16,000= $26,250

                   Remaining Life           

PAGE 15 GENERAL JOURNAL Score: 25/25 POST. REF. CREDIT DATE Dec. 31 ACCOUNT TITLE Depreciation Expense Accumulated Depreciati

Balland changes to the sum-of-the-years’-digits method.

Sum of the Years Digits Method

Year

Depreciation Base

Fraction

Depreciation Exp

Accumulated Depreciation at Year End

Book Value at Year End(cost-Accudepryear end)

5

226,000

5/21

60,000

The denominator of the fraction is the sum of the years of the asset's service life. Therefore, for an asset with a 6-year life, the sum is. The numerator of the fraction is the number of years remaining in the asset's life as of the beginning of the year.

Book Value at Beginning of Year 5 = 226,000

Sum of Years digits = n (n+1)/2 = 6(6+1)/2 = 42/2=21(10YEARS EST. LIFE – 4 YEARS AGO = 6)

Depreciation      =      (remaining book value – residual value) x sum of the years digits fraction

                   =       (226,000 – 16,000) x 6/21

=       210,000 x 6/21 = 60,000

PAGE 16 GENERAL JOURNAL Score: 25/25 DATE POST. REF. CREDIT Dec 31 ACCOUNT TITLE Depreciation Expense Accumulated Depreciatio

. Balland discovers that the estimated residual value has been ignored in the computation of depreciation expense

Depreciation without Residual Value = $366,000/10 years = $36,600/year

Incorrect Accumulated Depreciation = $36,600 x 4 years = $146,400

Correct Depreciation = 366,000 - 16,000/10 years = $35,000/year

Correct Accumulated Depreciation = $35,000 x 4 years = $140,000

Accumulated Depreciation = Incorrect Accumulated Depreciation - Correct Accumulated Depreciation

Accumulated Depreciation = 146,400 – 140,000 = 6,400      

Journal Entry

Accumulated Depreciation                $6,400

          Retained Earnings                                     $6,400       

Prior Period adjustment for error ($84,000 - $80,000)

Depreciation expense                        $35,000

Accumulated Depreciation                          $35,000

PAGE 16 GENERAL JOURNAL Score: 51/51 DATE POST. REF. CREDIT Dec 31 DEBIT 6,400.00 ACCOUNT TITLE Accumulated Depreciation Reta